A toy car of mass 0.620 kg is pressed against a spring. The spring constant is 250 N/m and the spring is compressed a distance of 0.031 m. How fast is the cart moving after it's released from the spring? Ignore the effects of friction.
Answer #1

In this problem, the energy stored in the energy is converted into the form of kinetic energy. i.e

2 2

muk

\Rightarrow v=\sqrt{\frac{kx^2}{m}}

Putting in the value

u-v250 x 6031- = 0.622 m/s

So, the speed of the car after its release is 0.622 m/s.